# Evaluating Intuitive Physics Understanding in Video Diffusion Models via Likelihood Preference

### v1

@@ Abstract
Intuitive physics understanding in video diffusion models plays an essential role in building general-purpose physically plausible world simulators, yet accurately evaluating such capacity remains a challenging task due to the difficulty in disentangling physics correctness from visual appearance in generation. To the end, we introduce LikePhys, a training-free method that evaluates intuitive physics in video diffusion models by distinguishing physically valid and impossible videos using the denoising objective as an ELBO-based likelihood surrogate on a curated dataset of valid-invalid pairs. By testing on our constructed benchmark of twelve scenarios spanning over four physics domains, we show that our evaluation metric, Plausibility Preference Error (PPE), demonstrates strong alignment with human preference, outperforming state-of-the-art evaluator baselines. We then systematically benchmark intuitive physics understanding in current video diffusion models. Our study further analyses how model design and inference settings affect intuitive physics understanding and highlights domain-specific capacity variations across physical laws. Empirical results show that, despite current models struggling with complex and chaotic dynamics, there is a clear trend of improvement in physics understanding as model capacity and inference settings scale.  [^1] Data and code are available at https://github.com/YuanJianhao508/LikePhys.
@@ 1 Introduction
Video diffusion models (VDMs) (Brooks et al., 2024; Google DeepMind, 2025; Zhou et al., 2024) have achieved impressive results in producing visually compelling videos, but they still often generate physically implausible outputs (Bansal et al., 2024; Motamed et al., 2025). Ensuring generative models learn the underlying physics that govern the dynamics of visual data is essential not only to improve the outputs’ quality (Kang et al., 2024; Li et al., 2025), but also to build realistic world models (LeCun, 2022; Ha & Schmidhuber, 2018) with applications in robotics (Hu et al., 2023b; Yang et al., 2023) and autonomous driving (Hu et al., 2023a; Zhao et al., 2025).
@@
However, it is challenging to evaluate how visual generative models learn and internalise physics. A classic line of work on general vision models uses the violation-of-expectation paradigm (Spelke, 1985; Baillargeon et al., 1985), which frames intuitive physics understanding as the ability to judge the plausibility of observed events (Riochet et al., 2018; Weihs et al., 2022), when presented with unrealistic videos obtained in simulation. However, extending this paradigm to generative models remains challenging. More recent approaches rely on Vision Language Models (VLMs) for plausibility judgments (Bansal et al., 2024; Guo et al., 2025) and text-conditioned compliance measures (Zhang et al., 2025; Meng et al., 2024a). Although these methods provide valuable insights, they usually fail to disentangle physics from visual appearance (Motamed et al., 2025) or introduce subjective biases (Wu & Aji, 2023) and interpretive variance (Gu et al., 2024) through VLM judgments, ultimately struggling to provide a grounded assessment of intuitive physics understanding.
@@
In this paper we propose an alternative evaluation method named LikePhys, that leverages the VDMs’ density estimation capacity (Li et al., 2023) rather than relying solely on generated outputs, as we show in Fig. 1. Inspired by the violation-of-expectation paradigm (Riochet et al., 2018), we start from the assumption that the capacity of a model to assign higher likelihood to physically-plausible visual sequences is correlated to its intuitive physics understanding. Then, we use a simulator to render paired videos. In one, we render physically-realistic phenomena, while in the other we introduce a controlled violation of physics. We keep the visual appearance consistent in the pair, ensuring that any difference resulting from processing the videos can be attributed solely to the breach of physics principles. Both rendered videos are then corrupted with noise and processed by the diffusion denoising network, where noise prediction losses serve as a proxy for sample likelihood (Ho et al., 2020). We then calculate the Plausibility Preference Error (PPE), i.e. we assume positive scores if the physically plausible sample has higher likelihood, and negative otherwise. Aggregating PPE results across pairs yields a single score that quantifies the intuitive physics understanding of a single model. We can compare scores across models to extract a ranking.
@@
To enable our evaluation, we construct a synthetic benchmark of twelve scenarios spanning Rigid Body Mechanics, Continuum Mechanics, Fluid Mechanics, and Optical Effects. Using this benchmark, we rank the performance in terms of PPE of 12 state-of-the-art VDMs with verification from human preference. Furthermore, we analyse the key factors in model design and inference settings influencing their intuitive physics understanding, and examine domain-specific capacity variations across the physics domain and laws to highlight the limitations of current VDMs. Empirical results demonstrate that our proposed metric is a robust proxy for physics understanding in VDMs, offering actionable insights into both their current limitations and their potential for progress. Our key contributions are the following:
@@ 2.1 Video Diffusion Models
VDMs have emerged as powerful generative frameworks for producing realistic video sequences (Esser et al., 2023; Brooks et al., 2024; Zhou et al., 2024; Girdhar et al., 2024). Early approaches such as VDM (Ho et al., 2022), AnimateDiff (Guo et al., 2024), and ModelScope (Wang et al., 2023) relied on 3D UNets (Ronneberger et al., 2015) to demonstrate the feasibility of spatio-temporal modeling. More recent systems, including CogVideoX (Yang et al., 2024), Hunyuan T2V (Kong et al., 2024), Wan (Wan et al., 2025), and LTX (HaCohen et al., 2024), adopt Transformer-based backbones in the Diffusion Transformer (DiT) style (Peebles & Xie, 2023; Vaswani et al., 2017), advancing long-sequence generation, visual quality, and inference efficiency. Despite these advances, it remains unclear whether current models capture underlying physical principles (Motamed et al., 2025). In this work, we propose a general evaluation protocol to assess the extent to which VDMs implicitly learn the laws of physics.
@@ 2.2 Intuitive Physics Understanding
Intuitive physics understanding is fundamental to a model’s ability to reason about and predict scenes under physics laws (Ates et al., 2020; Bear et al., 2021; Zhan et al., 2024; Meng et al., 2024b). One important line of work builds on the violation-of-expectation paradigm (Spelke, 1985; Baillargeon et al., 1985; Margoni et al., 2024), framing physics understanding in general vision models as the ability to judge the plausibility of observed events (Smith et al., 2019; Riochet et al., 2020; Weihs et al., 2022; Garrido et al., 2025). For example, IntPhys1/2 (Riochet et al., 2018; Bordes et al., 2025) assess intuitive physics using tightly controlled synthetic video pairs that differ only by a single violation of permanence, immutability, spatio-temporal continuity, or solidity to study each physics principle in isolation. When specifically applied to generative models, this protocol requires conditioned generation on context frames and computes the pixel reconstruction error on the target segment as plausibility proxy. However, it remains unclear how to adapt to text-conditioned VDMs that have no image-condition generation capacity. Moreover, these benchmarks are relatively constrained in rigid and solid object interactions. Broader coverage of deformable objects, fluids, and optical effects is limited.
@@ 2.3 Physics Evaluation in Video Generative Models
Some recent works focus on evaluating video generative models. A popular branch of methods (Zhang et al., 2025; Guo et al., 2025; Meng et al., 2024a) uses VLM with question–answering templates to assess adherence to physical laws in generated videos. For example, VideoPhy1/2 (Bansal et al., 2024; 2025) collect human-annotated synthetic videos and finetune VLM judges to align with human preferences on physical plausibility. However, these approaches are subject to visual appearance biases across different generative models (Wu & Aji, 2023) and interpretive variance across prompts and judging templates (Gu et al., 2024). Physics-IQ (Motamed et al., 2025) mitigates these issues through pixel-level pairwise comparisons with ground truth, but it relies on an image-conditioned architecture, and its applicability to text-to-video generation also remains unclear.
To move beyond these limitations, we take inspiration from the violation-of-expectation paradigm while removing the need for conditional generation or pixel-level alignment. Our method adapts the density estimation capability (Li et al., 2023) of VDMs and performs pairwise comparisons by directly evaluating sample likelihoods through the denoising loss proxy. This design keeps the metric model-agnostic and appearance-agnostic, avoiding confounds from visual artifacts while focusing directly on physics understanding.
@@ 3.1 Preliminaries on Video Diffusion Models
Diffusion Probabilistic Models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Nichol & Dhariwal, 2021) learn the data distribution by inverting a forward Markov noising process: $q({x}_{t}\mid{x}_{t-1})=\mathcal{N}\bigl({x}_{t};\,\sqrt{1-\beta_{t}}\,{x}_{t-1},\,\beta_{t}{I}\bigr),\quad\beta_{t}\in(0,1),$ A parameterised model then learns the reverse process $p_{\theta}({x}_{t-1}\mid{x}_{t})=\mathcal{N}\bigl({x}_{t-1};\,\mu_{\theta}({x}_{t},t),\,\sigma_{t}^{2}{I}\bigr)$ by minimising a noise prediction loss which serves as an ELBO-based surrogate for the negative log-likelihood. The noise prediction loss is calculated by reconstructing the ground truth noise $\epsilon$ injected on $x_{t}$ with a denoising network $\epsilon_{\theta}$, following:
@@ 3.2 Physics Understanding as Likelihood Preference
Our intuition is to formalise intuitive physics understanding in a distributional perspective. Let $p_{\rm phys}(x)$ denote the distribution over videos that strictly obey physical laws, where $x$ is a video sequence. We define its support as $\mathcal{M}_{\rm phys}\;=\;\bigl\{\,x\in\mathcal{X}:p_{\rm phys}(x)>0\bigr\}.$ By definition, any valid sample $x^{+}\in\mathcal{M}_{\rm phys}$ satisfies real-world physical laws, whereas any invalid sample $x^{-}\notin\mathcal{M}_{\rm phys}$ violates some of them. We consider a diffusion model $p_{\theta}$ has a perfect intuitive physics understanding if, for every valid–invalid pair $(x^{+},x^{-})$,
@@
Hence, if a diffusion model has learned the distribution correctly, it should assign higher likelihood to valid samples when they are paired with visually matched invalid samples. As shown in Eq. 1, the denoising loss is an ELBO-based likelihood surrogate. Thus, a lower denoising loss corresponds to a higher likelihood under $p_{\theta}$. For any valid–invalid pair $(x^{+},x^{-})$, we therefore have
$p_{\theta}(x^{+})\;>\;p_{\theta}(x^{-})\quad\Longleftrightarrow\quad\mathcal{L}_{\rm denoise}(\theta;x^{+})\;<\;\mathcal{L}_{\rm denoise}(\theta;x^{-}).$ (3)
To evaluate this likelihood-preference protocol in practice, we collect two sets of controlled video pairs following the violation-of-expectation paradigm (Margoni et al., 2024; Bordes et al., 2025):
$\mathcal{V}^{+}=\{x_{j}^{+}\}_{j=1}^{M},\qquad\mathcal{V}^{-}=\{x_{k}^{-}\}_{k=1}^{N},$ (4)
For each pair $(x^{+},x^{-})$, we perturb both with the same Gaussian noise at sampled diffusion timestep, pass them through the denoising network to predict the noise, and average the loss over timesteps as in Fig. 2. We say the model has violated expectation on the pair $(x_{j}^{+},x_{k}^{-})$ if it assigns no greater likelihood to the valid sample than to the invalid one. Aggregating over all $M\times N$ valid–invalid pairs, we define the Plausibility Preference Error (PPE):
$\text{PPE}\;=\;\frac{1}{M\,N}\sum_{j=1}^{M}\sum_{k=1}^{N}\mathbf{1}\Bigl[\mathcal{L}_{\rm denoise}(\theta;x_{j}^{+})\;\geq\;\mathcal{L}_{\rm denoise}(\theta;x_{k}^{-})\Bigr].$ (5)
A lower PPE indicates that $p_{\theta}$ consistently prefers valid samples. Thus, it measures the model’s likelihood preference for physically plausible videos, ultimately serving as a proxy for the learned intuitive physics.
@@ 3.3 Evaluation benchmark construction
To rigorously evaluate the learned physics, we require controlled video pairs that differ only in physical validity. In practice, it is difficult to obtain such matched pairs from natural data where physical laws cannot be systematically violated. We therefore construct a synthetic simulation benchmark spanning 12 physical scenarios across 4 domains, which allows precise control over physics parameters. All videos are rendered in Blender (Blender, 2018) at $512\times 512$ resolution with 60 frames. Each scenario is generated with 10 variations that vary both physics parameters and visual factors (e.g., object shape, texture, or environment). Within each variation, $N$ valid video strictly obeys the relevant law, while $M$ invalid variants introduce a single controlled violation (e.g., superelastic bounce or ghosted shadows). By holding camera angle, lighting, textures, and object geometry constant within a single variation, the valid and invalid videos differ only in physical plausibility, ensuring that any measured likelihood gap can be attributed to physics violations. As shown in Fig. 3, benchmark composition is as follows (see Apx. B for more details):
@@
4 Experiments
We evaluate the intuitive physics understanding of pre-trained VDMs using LikePhys and perform a systematic analysis to answer the following key questions: (1) Which models exhibit the strongest intuitive physics as measured by PPE? (2) How well does PPE align with human preference in physics correctness of generation? (3) To what extent does PPE disentangle visual appearance from intuitive physics? (4) What model design and inference factors influence models’ intuitive physics? (5) Across which physics domains and laws do current models fall short?
Settings: We benchmark pre-trained text-to-video diffusion models including AnimateDiff-SDv1.5/SDXL (Guo et al., 2024), ModelScope/ZeroScope (Wang et al., 2023), Mochi (Team, 2024), CogVideoX-2B/5B (Yang et al., 2024), Wan2.1-1.3B/14B (Wan et al., 2025), Hunyuan T2V (Kong et al., 2024) and LTX (HaCohen et al., 2024). To maximise performance, we adopt each model’s recommended inference settings from its official repository and resample the frame rate and spatial resolution of generated video samples to match. We then perform zero‐shot evaluation using the proposed LikePhys protocol, measuring performance via the PPE metric defined in Eq. 5 and ranking models. For each physics scenario, we report the average error over all valid–invalid pairs. For the likelihood estimation, we use the same DDIM Euler (Song et al., 2020) scheduler with 10 timesteps uniformly sampled through the noise scale. We use a fixed prompt template for each physics scenario. More details of the evaluation implementation (Apx. A), evaluation dataset construction and specification (Apx. B), inference settings (Apx. C), experiment settings (Apx. D), additional ablation study on design factors of evaluation method (Apx. E), further applicability to other intuitive physics benchmarks (Apx. F) and visual samples (Apx. G) are provided in the appendix.
4.1 Main results
@@
Model ranking
In Tab. 1, we benchmark twelve VDMs and order them by increasing PPE. The results reveal large performance gaps across architectures. Early UNet-based models (Ronneberger et al., 2015) such as AnimateDiff and ZeroScope often record error rates above 50%, reflecting limited ability to distinguish physically valid videos from invalid ones. In contrast, more recent DiT-based designs (Peebles & Xie, 2023), including the top three performers Hunyuan T2V (43.6%), Wan2.1–T2V–14B (43.8%), and CogVideoX1.5–5B (43.8%), achieve substantially lower errors across multiple scenarios (see Sec. 4.2 for further discussion).
Moreover, while there is a clear trend of improvement in recent architectures, only a handful of models significantly achieve PPE lower than the 50% random-guess threshold in more than a few scenarios, potentially due to simplicity biases, where physically implausible interactions appear visually easier to model than valid ones (Shah et al., 2020; Geirhos et al., 2020). These results highlight that while modern VDMs start to internalise physical principles, their performance varies substantially across different physics domains (see Sec. 4.3 for further analysis), indicating significant room for progress in physically-accurate training of video generators.
Alignment to Human Preference

### v2

# LikePhys: Evaluating Intuitive Physics Understanding in Video Diffusion Models via Likelihood Preference

@@ Abstract
Intuitive physics understanding in video diffusion models plays an essential role in building general-purpose physically plausible world simulators, yet accurately evaluating such capacity remains a challenging task due to the difficulty in disentangling physics correctness from visual appearance in generation. To the end, we introduce LikePhys, a training-free method that evaluates intuitive physics in video diffusion models by distinguishing physically valid and impossible videos using the denoising objective as an ELBO-based likelihood surrogate on a curated dataset of valid-invalid pairs. By testing on our constructed benchmark of twelve scenarios spanning over four physics domains, we show that our evaluation metric, Plausibility Preference Error (PPE), demonstrates strong alignment with human preference, outperforming state-of-the-art evaluator baselines. We then systematically benchmark intuitive physics understanding in current video diffusion models. Our study further analyses how model design and inference settings affect intuitive physics understanding and highlights domain-specific capacity variations across physical laws. Empirical results show that, despite current models struggling with complex and chaotic dynamics, there is a clear trend of improvement in physics understanding as model capacity and inference settings scale. Our data and code will be open source.
@@ 1 Introduction
Video diffusion models (VDMs) (sora; deepmind_veo3; polyak2024movie) have achieved impressive results in producing visually compelling videos, but they still often generate physically implausible outputs (bansal2024videophy; motamed2025physicsiq). Ensuring generative models learn the underlying physics that govern the dynamics of visual data is essential not only to improve the outputs’ quality (kang2024far; li2025worldmodelbench), but also a essential pre-requisit for them to serve as reliable world models (lecun2022path; ha2018world) with applications in robotics (yang2023learning) and autonomous driving (hu2023gaia; zhao2025drivedreamer).
@@
However, it is challenging to evaluate how visual generative models learn and internalise physics. A classic line of work on general vision models uses the violation-of-expectation paradigm (spelke1985preferential; baillargeon1985object), which frames intuitive physics understanding as the ability to judge the plausibility of observed events (riochet2018intphys; WeihsEtAl2022InfLevel), when presented with unrealistic videos obtained in simulation. However, extending this paradigm to generative models remains challenging. More recent approaches rely on Vision Language Models (VLMs) for plausibility judgments (bansal2024videophy; guo2025t2vphysbench) and text-conditioned compliance measures (meng2024towards). Although these methods provide valuable insights, they usually fail to disentangle physics from visual appearance (motamed2025physicsiq) or introduce subjective biases (wu2023style) and interpretive variance (gu2024llmjudge) through VLM judgments, ultimately struggling to provide a grounded assessment of intuitive physics understanding.
@@
In this paper we propose an alternative evaluation method named LikePhys, that leverages the VDMs’ density estimation capacity (li2023diffusion) rather than relying solely on generated outputs, as we show in Fig. 1. Inspired by the violation-of-expectation paradigm (riochet2018intphys), we start from the assumption that the capacity of a model to assign higher likelihood to physically-plausible visual sequences is correlated to its intuitive physics understanding. Then, we use a simulator to render paired videos. In one, we render physically-realistic phenomena, while in the other we introduce a controlled violation of physics. Importantly, we keep the visual appearance consistent in the pair, ensuring that any difference resulting from processing the videos can be attributed solely to the breach of physics principles. Both rendered videos are then corrupted with noise and processed by the diffusion denoising network, where noise prediction losses serve as a proxy for sample likelihood (ho2020denoising). We then calculate the Plausibility Preference Error (PPE), i.e. we assume positive scores if the physically plausible sample has higher likelihood, and negative otherwise. Aggregating PPE results across pairs yields a single score that quantifies the intuitive physics understanding of a single model.
@@
To enable our evaluation, we construct a synthetic benchmark of twelve scenarios spanning Rigid Body Mechanics, Continuum Mechanics, Fluid Mechanics, and Optical Effects. Importantly, we design each physics scenario to be relatively simple, with clearly attributable governing physics dynamics, and consistent in appearance, so that any likelihood variation due to model preferences for visual quality cancels out during pairwise preference comparisons. Using this benchmark, we rank the performance of 12 state-of-the-art VDMs in terms of PPE, based on human preference-based verification. Furthermore, we analyse the key factors in model design and inference settings influencing their intuitive physics understanding, and examine domain-specific capacity variations across the physics domain and laws to highlight the limitations of current VDMs. Empirical results demonstrate that our proposed metric is a robust proxy for physics understanding in VDMs, offering actionable insights into both their current limitations and their potential for progress. Our key contributions are the following:
@@ 2.1 Video Diffusion Models
VDMs have emerged as powerful generative frameworks for producing realistic video sequences (esser2023structure; sora; zhou2024allegro; girdhar2024EMU). Early approaches such as VDM (ho2022video), AnimateDiff (guo2023animatediff), and ModelScope (wang2023modelscope) relied on 3D UNets (ronneberger2015u) to demonstrate the feasibility of spatio-temporal modeling. More recent systems, including CogVideoX (yang2024cogvideox), Hunyuan T2V (kong2024hunyuanvideo), Wan (wan2025), and LTX (HaCohen2024LTXVideo), adopt Transformer-based backbones in the Diffusion Transformer (DiT) style (peebles2023dit; vaswani2017attention), advancing long-sequence generation, visual quality, and inference efficiency. Despite these advances, it remains unclear whether current models capture underlying physical principles (motamed2025physicsiq). In this work, we propose a general evaluation protocol to assess the extent to which VDMs implicitly learn the laws of physics.
@@ 2.2 Intuitive Physics Understanding
Intuitive physics understanding is fundamental to a model’s ability to reason about and predict scenes under physics laws (ates2020craft; bear2021physion; guanqi; meng2024phybench). One important line of work builds on the violation-of-expectation paradigm (spelke1985preferential; baillargeon1985object; margoni2024violation), framing physics understanding in general vision models as the ability to judge the plausibility of observed events (smith2019modeling; riochet2020occlusion; WeihsEtAl2022InfLevel; garrido2025intuitive). For example, IntPhys1/2 (riochet2018intphys; bordes2025intphys) assess intuitive physics using tightly controlled synthetic video pairs that differ only by a single violation of permanence, immutability, spatio-temporal continuity, or solidity to study each physics principle in isolation. These works are mostly built on simulated data and have been successful in evaluating the physics understanding of various vision models (garrido2025intuitive; bordes2025intphys). However, extending this protocol to visual generative models is non-trivial: it typically relies on context-conditioned generation and pixel reconstruction on target frames as a plausibility proxy, and it remains unclear how to adapt this setup to text-conditioned video diffusion models that lack image conditioning and are not naturally discriminative.
@@ 2.3 Physics Evaluation in Video Generative Models
Some recent works focus on evaluating video generative models. A popular branch of methods (guo2025t2vphysbench; meng2024towards) uses VLM with question–answering templates to assess adherence to physical laws in generated videos. For example, VideoPhy1/2 (bansal2024videophy; bansal2025videophy2) collect human-annotated synthetic videos and finetune VLM judges to align with human preferences on physical plausibility. However, these approaches are subject to visual appearance biases across different generative models (wu2023style) and interpretive variance across prompts and judging templates (gu2024llmjudge). Physics-IQ (motamed2025physicsiq) and Morpheus (zhang2025morpheus) address these issues by comparing generated videos with paired real recordings via pixel/object-mask analyses and by fitting dynamical models to extract physical quantities, then scoring adherence to physical conservation laws, respectively. However, they rely on image-conditioned generation, and its extension to text-to-video generation also remains unclear.
To move beyond these limitations and provide an alternative evaluation from a likelihood preference perspective, we take inspiration from the violation-of-expectation paradigm while removing the need for conditional generation or pixel-level alignment. Our method adapts the density estimation capability (li2023diffusion) of VDMs and performs pairwise comparisons by directly evaluating sample likelihoods through the denoising loss proxy. This design keeps the metric model-agnostic and appearance-agnostic, avoiding confounds from visual artifacts while focusing directly on physics understanding.
@@ 3.1 Preliminaries on Video Diffusion Models
Diffusion Probabilistic Models (sohl2015deep; ho2020denoising; nichol2021improved) learn the data distribution by inverting a forward Markov noising process: $q({x}_{t}\mid{x}_{t-1})=\mathcal{N}\bigl({x}_{t};\,\sqrt{1-\beta_{t}}\,{x}_{t-1},\,\beta_{t}{I}\bigr),\quad\beta_{t}\in(0,1),$ A parameterised model then learns the reverse process $p_{\theta}({x}_{t-1}\mid{x}_{t})=\mathcal{N}\bigl({x}_{t-1};\,\mu_{\theta}({x}_{t},t),\,\sigma_{t}^{2}{I}\bigr)$ by minimising a noise prediction loss which serves as an ELBO-based surrogate for the negative log-likelihood. The noise prediction loss is calculated by reconstructing the ground truth noise $\epsilon$ injected on $x_{t}$ with a denoising network $\epsilon_{\theta}$, following:
@@ 3.2 Physics Understanding as Likelihood Preference
Our intuition is to formalise intuitive physics understanding in a distributional perspective. Let $p_{\rm phys}(x)$ denote the distribution over videos that strictly obey physical laws, where $x$ is a video sequence. We define its support as $\mathcal{M}_{\rm phys}\;=\;\bigl\{\,x\in\mathcal{X}:p_{\rm phys}(x)>0\bigr\}.$ By definition, any valid sample $x^{+}\in\mathcal{M}_{\rm phys}$ satisfies real-world physical laws, whereas any invalid sample $x^{-}\notin\mathcal{M}_{\rm phys}$ violates some of them. We consider a diffusion model $p_{\theta}$ that learn has a perfect intuitive physics understanding if, for every valid–invalid pair $(x^{+},x^{-})$,
@@
Hence, if a diffusion model has learned the distribution correctly, it should assign a higher likelihood to valid samples when they are paired with visually matched invalid samples. As shown in Eq. 1, the denoising loss is an ELBO-based likelihood surrogate. Thus, a lower denoising loss corresponds to a higher likelihood under $p_{\theta}$. For any valid–invalid pair $(x^{+},x^{-})$, we therefore have
$p_{\theta}(x^{+})>p_{\theta}(x^{-})\Longleftrightarrow\mathcal{L}_{\rm denoise}(\theta;x^{+})<\mathcal{L}_{\rm denoise}(\theta;x^{-}).$ (3)
To evaluate this likelihood–preference protocol in practice, we first identify 12 physics scenarios. For each scenario, we create $R$ controlled variations to account for variability in both physics parameters and confounding visual factors (with $R=10$ in our benchmark; see Sec. 3.3). For each variation $r\in\{1,\dots,R\}$, we collect sets of controlled video pairs following the violation-of-expectation paradigm (margoni2024violation; bordes2025intphys):
$\mathcal{V}_{r}^{+}=\{x_{r,j}^{+}\}_{j=1}^{M_{r}},\qquad\mathcal{V}_{r}^{-}=\{x_{r,k}^{-}\}_{k=1}^{N_{r}},$ (4)
where $\mathcal{V}_{r}^{+}$ and $\mathcal{V}_{r}^{-}$ denote valid and invalid samples for variation $r$, respectively. For all video samples in $\mathcal{V}_{r}^{+}$ and $\mathcal{V}_{r}^{-}$, we ensure they are simple and consistent in visual appearance, with the only differences arising from violations of the governing physics dynamics. This ensures that any likelihood variation due to model preferences on visual style cancels out, enabling an unbiased comparison across different video diffusion models.
We then perturb both videos in each pair with the same Gaussian noise at sampled diffusion timesteps, pass them through the denoising network to predict the noise, and average the loss over timesteps as in Fig. 2, obtaining a denoising loss $\mathcal{L}_{\mathrm{denoise}}(\theta;x)$ for each video. For a given variation $r$, we say the model has violated expectation on the pair $(x_{r,j}^{+},x_{r,k}^{-})$ if it assigns no greater likelihood (i.e., no lower denoising loss) to the valid sample than to the invalid one. Aggregating over all $N\times M$ pairs within each variation and then averaging across variations, we define the Plausibility Preference Error (PPE):
$\text{PPE}=\frac{1}{R}\sum_{r=1}^{R}\frac{1}{M_{r}\,N_{r}}\sum_{j=1}^{M_{r}}\sum_{k=1}^{N_{r}}\mathbf{1}\Bigl[\mathcal{L}_{\mathrm{denoise}}(\theta;x_{r,j}^{+})\geq\mathcal{L}_{\mathrm{denoise}}(\theta;x_{r,k}^{-})\Bigr],$ (5)
where $\mathbf{1}[\cdot]$ is the indicator function and the inner average corresponds to the rate that a model mis-assign higher likelihood to invalid video sample for each variation $r$. A lower PPE indicates that $p_{\theta}$ consistently prefers valid samples across variations, thus measuring the model’s likelihood preference for physically plausible videos and serving as a proxy for its learned intuitive physics.
@@ 3.3 Evaluation Benchmark Construction
To rigorously evaluate the learned physics, we require controlled video pairs that differ only in physical validity. In practice, it is infeasible to obtain such matched pairs from real-world data, where physics laws cannot be systematically violated. We therefore construct a synthetic simulation benchmark spanning 12 physical scenarios across 4 domains, which allows precise control over physics parameters. All videos are rendered in Blender (blender) at $512\times 512$ resolution with 60 frames. Each scenario is generated with $R=10$ variations that vary both physics parameters and visual factors (e.g., object shape, texture, or environment). Within each variation, we generate $M$ valid videos that strictly obey the relevant law, and $N$ invalid variants that introduce a single controlled violation (e.g., superelastic bounces or ghosted shadows). By holding camera angle, lighting, textures, and object geometry constant within a single variation, the valid and invalid videos differ only in physical plausibility, ensuring that any measured likelihood gap can be attributed to physics violations. As shown in Fig. 3, the benchmark composition is as follows (see Apx. B for more details):
@@
4 Experiments
We evaluate the intuitive physics understanding of pre-trained VDMs using LikePhys and perform a systematic analysis to answer the following key questions: (1) Which models exhibit the strongest intuitive physics as measured by PPE? (2) How well does PPE align with human preference in physics correctness of generation? (3) To what extent does PPE disentangle visual appearance from intuitive physics? (4) What model design and inference factors influence models’ intuitive physics? (5) Across which physics domains and laws do current models fall short?Settings: We benchmark pre-trained text-to-video diffusion models including AnimateDiff-SDv1.5/SDXL (guo2023animatediff), ModelScope/ZeroScope (wang2023modelscope), Mochi (genmo2024mochi), CogVideoX-2B/5B (yang2024cogvideox), Wan2.1-1.3B/14B (wan2025), Hunyuan T2V (kong2024hunyuanvideo) and LTX (HaCohen2024LTXVideo). To maximise performance, we adopt each model’s recommended inference settings from its official repository and resample the frame rate and spatial resolution of generated video samples to match. We then perform zero‐shot evaluation using the proposed LikePhys protocol, measuring performance via the PPE metric defined in Eq. 5 and ranking models. For each physics scenario, we report the average error over all valid–invalid pairs. For the likelihood estimation, we use the same DDIM (song2020denoising) scheduler with 10 timesteps uniformly sampled through the noise scale. We use a fixed prompt template for each physics scenario. More details of the evaluation implementation (Apx. A), evaluation dataset construction and specification (Apx. B), inference settings (Apx. C), experiment settings (Apx. D), additional ablation study on design factors of evaluation method (Apx. E), further applicability to other intuitive physics benchmarks (Apx. F) and visual samples (Apx. G) are provided in the appendix.
4.1 Main results